rational equations: Solve for X using LCD method
x+1/3 + x-3/4 = 1/6

Answer by edjones(8007)

x + 1/3 + x - 3/4 = 1/6
12x+4+12x-9=2 LCD=12
24x-5=2
24x=7
x=7/24
.
Ed